OpenStudy (anonymous):

-3x = 5 - y 2y = 6x + 10 How do I solve these systems of equations by graphing?

OpenStudy (anonymous):

In order to graph I will simplify using substitution

OpenStudy (anonymous):

-3x=5-y +5 +5 _____________________ -3x+5=-y 3x-5=y

OpenStudy (anonymous):

I don't understand what you did with the five in the first line, isn't the correct thing to do is subtract?

OpenStudy (anonymous):

I added 5 on both sides

OpenStudy (anonymous):

Since 5 and -3x aren't like terms we leave them as is